# MedicoSAM: Towards foundation models for medical image segmentation

### v1

@@ 2.3 Interactive 3D Segmentation
Unlike SAM2, which supports image and video segmentation, SAM supports segmentation only for 2D data. Follow-up work has implemented interactive segmentation for videos or volumetric data in medical images, e.g. for CT [lei2023medlsam] and Bone CT [maggtraining]. Here, we use the implementation from [archit2023segment], which is based on prompt propagation. Briefly, a user annotates an object in one or multiple slices with prompts and SAM is applied to these slices. Then, the segmentation mask(s) are projected to adjacent slices, prompts are derived from them, and segmentation is run for these prompts. The process is repeated until the object is segmented throughout the whole volume or a stopping criterion based on the IOU between adjacent slices is met. Multiple options are provided for deriving prompts from projected masks: using a single positive point prompt placed at the mask’s center, using multiple point prompts derive from the mask, using the bounding box derived from the mask, using the bounding box and low-resolution version of the mask, and combinations of these options. A user can correct the segmentation by annotating slices with manual prompts and rerunning the segmentation.
@@ 2.4 Semantic Segmentation
On its own, SAM can only be used for interactive segmentation. In [kirillov2023segment] a method for automatic instance segmentation, called automatic mask generation, is proposed. However, this method only supports automatic instance segmentation and not semantic segmentation, which is the more relevant task for medical images. Conceptually, SAM does not learn explicit semantic knowledge, since it is only trained to distinguish objects without class information. Most approaches for semantic segmentation in medical images build on the U-Net [ronneberger2015u, cciccek20163d], nnU-NET [isensee2021nnu] being the most popular one. To evaluate different pretrained SAM models for semantic segmentation, we finetune them for specific tasks using annotated data. We implement an architecture for 2D segmentation and one for 3D segmentation, see also Fig. 2b.
For 2D segmentation we implement SAMed [zhang2023customized]. It updates the SAM architecture by replacing input prompts with a learnable default prompt and changing the number of heads in the multi-mask decoder branch from three to the number of classes. The resulting model can be trained for semantic segmentation by updating parameters based on a loss between labels and predictions. We use the combination of Cross Entropy and Dice loss. Note that SAMed uses LoRA [hu2021lora] in the image encoder. We compare two approaches, updating all parameters of the image encoder or using LoRA.
For 3D segmentation we implement MA-SAM [chen2023ma]. This approach follows a similar design as SAMed, introducing additional changes to the image encoder. It flattens batch and depth dimension so that all image patches are processed by the transformer blocks. To make use of depth information it introduces 3D adapter layers. These layers decrease the number of features per token, rearrange tokens into a volumetric representation, apply a 3x1x1 convolution, project the number of features back and flatten the batch and depth axis. Each transformer layer is augmented with two of these adapters, one before the attention layer and the other after it. The parameters of the 3D adapters are randomly initialized and updated during training. The other parameters in the attention layers are either all updated or updated through LoRA (same as for 2D).
@@ 2.5 Evaluation
For interactive 2D segmentation we adopt the evaluation procedure of [archit2023segment]. This approach simulates iterative user-based annotation. It requires object mask annotations. For a given object, a single prompt is sampled, either a point or a box. The object is then iteratively corrected by sampling point prompts from errors in the prediction. In each iteration a positive point prompt is sampled from the region where the prediction is missing (prediction is negative, annotation is positive) and a negative point prompt is sampled from the region where the prediction should not be (prediction is positive, annotation is negative). The Dice coefficient between true mask and prediction is computed for the initial segmentation and each correction iteration. For interactive 3D segmentation we evaluate the initial segmentation derived from a point prompt (randomly sampled from the object in the central slice) and from a box prompt (also in the central slice). We do not simulate iterative correction of the masks due to the higher computational demand of 3D segmentation. We run a grid search over the different options for deriving prompts, see Sec. 2.3, on separate validation data.
@@ 3 Experiments
We finetune three models based on SAM ViT-b, initialized with the weights from [kirillov2023segment]: MedSAM*, SimpleFT* and MedicoSAM, see Sec. 2.2 for an explanation of the three different training set-ups. The three models are trained on the publicly available subset of SA-Med2D-20M [ye2023sa], which contains 3,7 million images and 15,8 million masks. We use 50% of this dataset for training and 10% for validation, with a stratified split over the different modalities in the dataset. The remaining 40% of the dataset are not used to lower the training time per epoch. The models are trained on 8 A100 GPUs with 80GB VRAM for 300,000 iterations with a batch size of 7 per GPU, corresponding to 9 epochs. We use the AdamW optimizer [loshchilov2017decoupled] with an initial learning rate of $55e-$ 5 and a scheduler that reduces the learning rate by a factor of 0.9 after each epoch.
We evaluate these models and compare to the original SAM model and already published models for medical images for five different applications: interactive segmentation (2D and 3D), finetuning for semantic segmentation (2D and 3D) and integration with user-friendly tools.
Figure 3: a) Overall results for interactive 2D segmentation. We report the Dice coefficient for simulated interactive segmentation. Each bar corresponds to the result of a correction iteration, starting either from a point (green) or a box (yellow) prompt. The result after correction is highlighted in dark green / dark purple. We compare 8 different models. Models that were trained by us are marked with a * or in bold for the MedicoSAM model. b) Interactive segmentation results for 16 individual datasets. We report the absolute difference of the Dice coefficient compared to the original SAM and report only the results for the initial and final segmentation.
@@ 3.1 Interactive Segmentation
We evaluate eight different models for interactive 2D segmentation, three we have finetuned and five published ones (including the original SAM and SAM2), see Sec. 2.2 for details. We use 16 external datasets that are not part of our training dataset to evaluate generalization capabilities. These datasets represent a variety of medical segmentation tasks from CT [pepe2024segmentation, bolelli2023tooth, podobnik2023han], dermoscopy [glaister2014automatic], endoscopy [sanchez2020piccolo, maqbool2020m2caiseg], MRI [macdonald2023duke, van2024lumbar], ophthalmology [kovalyk2022papila], ultrasound [leclerc2019deep, lu2022jnu, jiang2024microsegnet], and X-Ray [jaeger2013automatic, lee2017curated, cervantes2019automatic, siiimacr]. Here, 3D dataset are split into separate images.
Figure 4: Results for interactive 3D segmentation for 6 different datasets. We report the difference in Dice score compared to SAM for four other models. Segmentations are derived from a single point (green) or box (yellow) prompt placed in the central slice for each object in the respective dataset. We use the implementation of [archit2023segment] for methods using SAM, determining the best method for prompt propagation on a separate validation set, see also Sec. 2.3. SAM2 supports 3D segmentations by default.
@@
The results averaged over all datasets are shown in Fig. 3 a). We report the results for initial prompt-based segmentation and correction iterations, see also Sec. 2.5, using a point or box as initial prompt. Fig. 3 b) shows the results for all individual datasets, where we only report the results for the initial and final segmentation, corresponding to the last correction iteration. We report the difference in the Dice coefficient compared to SAM. Here, MedicoSAM is the only model that clearly improves upon SAM for all settings. The other finetuned models either only improve for the initial segmentation (MedSAM*, SimpleFT*) or lead to an overall worse segmentation performance. SAM2 performs slightly worse compared to (original) SAM. A qualitative comparison of model predictions is shown in Fig. 7 (supp. material).
@@
We evaluate interactive 3D segmentation for the initial segmentation derived from a point prompot or a box prompt for 6 different external datasets from MRI [cancer2015comprehensive, macdonald2023duke], CT [heller2021state, osic-pulmonary-fibrosis-progression], and ultrasound [jiang2024microsegnet, kronke2022tracked]. We compare five different models. For the models based on SAM, we use the method described in Sec. 2.3 and we find the best setting with a grid search on a separate validation set. SAM2 supports interactive 3D segmentation as is (by interpreting the 3D data as a video), and we do not perform a grid search to optimize parameters for inference. The overall results are shown in Fig. 1 c) (except for SimpleFT) with results for individual datasets in Fig. 4. Only MedicoSAM improves consistently. The fact that SAM2 is used as is, without optimizing parameters in a grid search, may disfavor it. A qualitative comparison of model predictions is shown in Fig. 8 (supp. material).
Figure 5: Semantic segmentation results. We train an architecture for 2D and 3D segmentation. The names in italic font indicate 2D data and in bold font 3D data. We compare different pretrained models, original SAM, MedSAM, SimpleFT SAM, and MedicoSAM. We also compare updating all parameters of the encoder with LoRA . The three best methods are colored, in case MedicoSAM is not among them it is marked with a dashed outline. The red line marks the nnU-Net result.
@@ 3.2 Semantic Segmentation
We evaluate the use of different pretrained SAM models for semantic segmentation in 2D and 3D, using the implementations described in Sec. 2.4. We compare updating all parameters in the image encoder with parameter updates using LoRA [hu2021lora]. For 2D segmentation we use 6 external datasets from dermoscopy [codella2019skin], mammography [lee2017curated], narrow band imaging [sanchez2020piccolo], optical coherence tomography [ye2023oimhs], panoramic radiographs [roman2021panoramic], and X-Ray angiography [cervantes2019automatic]. For 3D segmentation we use 6 external datasets from CT [osic-pulmonary-fibrosis-progression], MRI [macdonald2023duke, marcus2007open, cancer2015comprehensive], and ultrasound [jiang2024microsegnet, kronke2022tracked]. We use separate splits for training and evaluation. The results are shown in Fig. 5. We also report the results for nnU-Net [isensee2021nnu] trained on the training split of the respective dataset, using the default nnU-Net v2 setup.
Here, we do not see an advantage in using a domain-specific pretrained backbone (MedSAM, SimpleFT, MedicoSAM) over the initial SAM model. The architectures initialized with SAM weights perform on par or slightly better than initialization with these models. Furthermore, we don’t see a clear advantage in the use of SAM-based semantic segmentation compared to nnU-Net. While their performance is better overall in 2D, due to the poor performance of nnU-Net on two of the datasets, nnU-Net is the best method for 3D segmentation. Here, we have to stress that this comparison disfavors nnU-Net by design, as it is always trained from scratch, whereas the other methods are initialized with pretrained encoders. Finally, we find that results with LoRA are always slightly worse compared to updating all parameters in the image encoder.
@@ 3.3 Tool Integration
An important practical aspect in improving SAM as a foundation model for medical images is the integration with graphical tools for data annotation. Hence, models should not introduce changes to the SAM architecture that make the model incompatible with the SAM library and tools using it. We check this for three models, MedSAM, SAM-Med2D and MedicoSAM, and four different tools: two napari plugins [napari-sam, archit2023segment] and two 3D Slicer extensions [liu2023samm, yildiz2024segmentwithsam]. Tab. 1 shows the compatibility. We find that SAM-Med2D does not work in any of the tools because it uses adapters and changes the image input size. MedSAM and MedicoSAM work in all of the tools with at most small code changes. We qualitatively compare the models for data annotation with these tools in Fig. 6.
@@
Figure 6: Using SAM, MedSAM and MedicoSAM in four different tools for interactive segmentation. The top row shows the reference image and the object that is annotated. The four rows below show interactive segmentation results with the different tools. Data on the left hand side shows a section from a abdominal MRI scan [ji2022amos], the right hand side shows a section from a cervical CT scan [landman2015miccai].
4 Conclusion
We have presented a comprehensive study of SAM for medical images, by evaluating the impact of different finetuning objectives on interactive segmentation and semantic segmentation. We found that interactive segmentation can improve clearly due to domain specific finetuning, compared both to original SAM and SAM2. In this context, the choice of the objective is crucial, and only the complex objective used in SAM preserves interactive segmentation capabilities, whereas simpler objectives as used in in [ma2024segment, gu2024build] lead to diminishing capabilities for iterative interactive segmentation. For semantic segmentation we find that using a domain specific backbone does not provide a clear benefit, at least using the approaches we study. Further, we don’t see a clear advantage of SAM-based architectures over nnU-Net [isensee2021nnu]. Finally, we show that models based on SAM should adhere to the original architecture to enable integration within existing tools. We believe that this fact is especially important since, according to our findings, their main advantage is improved interactive data annotation, which relies on such tools.
Our contribution provides important updates to prior work on SAM for medical images. In [ma2024segment], the authors claim that their model, MedSAM, improves interactive segmentation when using a simple finetuning objective using only box prompts. However, according to our results, this only holds true for annotation from bounding boxes, and the model loses the capability for correction with additional prompts. Given that this capability is one of the main strengths of SAM, we believe that this leads to a worse model for practical applications; [ma2024segment] misses this observation due to a simplistic evaluation procedure that only evaluates box prompts. While SAM-Med2D [cheng2023sam] uses a comprehensive finetuning objective and evaluation procedure, it uses adapters and changes the input image size, making it incompatible with existing tools. In [gu2024build], the authors study SAM as basis for semantic segmentation. In addition to finetuning with a prompt-based objective, they also study self-supervised finetuning. They find that prompt-based finetuning does not improve the performance for downstream semantic segmentation, which is consistent with our results, and that self-supervised finetuning may be slightly beneficial (we do not study this setting, because it will very likely lead to catastrophic forgetting of interactive segmentation capabilities). Furthermore, they claim that finetuing SAM shows advantages over nnU-Net, and that LoRA-based updates are superior to full parameter updates. Both claims are contrary to our results and may be explained due to different choices in training parameters, including for the nnU-Net baselines, and evaluation datasets. They also study data-efficient finetuining on only 5 training images, which we did not investigate.
We believe that our empirical results will provide valuable context for future efforts to adapt more powerful vision foundation models, e.g. SAM2, to medical images, marking an important step towards building a universal model for medical image segmentation. Besides these empirical findings, we publish our best model, MedicoSAM. It can be easily integrated within existing tools for data annotation. According to our results, it is the best currently available model for interactive segmentation of medical data, and we thus believe that it will provide great practical value.

### v2

@@ 2.3 Interactive 3D Segmentation
Unlike SAM2, which supports image and video segmentation, SAM can only segment 2D images. Follow-up work has implemented interactive segmentation for videos or volumetric data in medical images, e.g. for CT [41, 52]. Here, we use the implementation from [1], which is based on prompt propagation. Briefly, SAM segments an object in one or multiple slices based on given prompts. Then, the segmentation mask(s) are projected to adjacent slices, prompts are derived from them, and segmentation is run for these prompts. The process is repeated until the object is segmented throughout the whole volume or a stopping criterion based on the IOU between adjacent slices is met. Multiple options are provided for deriving prompts from projected masks: using a single positive point prompt placed at the mask’s center, using multiple point prompts derive from the mask, using the bounding box derived from the mask, using the bounding box and low-resolution version of the mask, and combinations of these options. A user can correct the segmentation by annotating slices with manual prompts and rerunning the segmentation.
@@ 2.4 Semantic Segmentation
SAM itself can only be used for interactive segmentation. In [33] a method for automatic instance segmentation, called automatic mask generation, is proposed. However, it does not support semantic segmentation, which is more relevant for medical images. Conceptually, SAM does not learn explicit semantic knowledge, since it is only trained to distinguish individual objects in an image. To evaluate different pretrained SAM models for semantic segmentation, we further finetune them for specific tasks using annotated data, based on architectural changes for 2D and 3D segmentation.
For 2D segmentation we add a UNETR-like [20] convolutional segmentation decoder. It processes and upsamples the embeddings predicted by the image encoder, in this case the pretrained ViT-b encoder, to predict a semantic segmentation. Prompt encoder and original mask encoder of SAM are discarded. This model is trained for semantic segmentation with a loss between predictions and semantic labels. Here, we use a combination of Cross Entropy and Dice loss. We also study a variant where the segmentation decoder is pretrained, see Sec. 2.2 for details. The segmentation decoder is marked in orange in Fig. 2a. The choice of this architecture for semantic segmentation was motivated by similar approaches for automatic segmentation with SAM in [1, 71].
For 3D segmentation we adopt the implementation of MA-SAM [8] to extend the image encoder to volumetric data. This is achieved by flattening the batch and depth dimensions so that the image patches extracted from an input volume can all be processed by the encoder. To make use of depth information, additional adapter layers are introduced. These layers decrease the number of features per token, rearrange tokens into a volumetric representation, apply a 3x1x1 convolution, project the number of features back, and flatten the batch and depth axis. Each transformer layer is augmented with two of these adapters, one before and one after the attention layer. The parameters of the adapters are randomly initialized. Unlike MA-SAM, which re-uses SAM’s mask decoder, we use our new 2D segmentation decoder. We apply it slice-by-slice to the image embeddings, then stack its outputs and apply a 3D convolution layer to predict a volumetric segmentation. This design enables to also initialize the pretrained decoder weights. The architecture is shown in Fig. 2b.
@@ 2.5 Evaluation
For interactive 2D segmentation we adopt the evaluation procedure of [1]. This approach simulates iterative user-based annotation. It requires object mask annotations. For a given object, a single prompt is sampled, either a point or a box. The object is then iteratively corrected by sampling point prompts from errors in the prediction. In each iteration a positive point prompt is sampled from the region where the prediction is missing (prediction is negative, annotation is positive) and a negative point prompt is sampled from the region where the prediction should not be (prediction is positive, annotation is negative). The Dice coefficient between true mask and prediction is computed for the initial segmentation and each correction iteration. For interactive 3D segmentation we evaluate the initial segmentation derived from a point prompt (randomly sampled from the object in the central slice) and from a box prompt (also in the central slice). We do not simulate iterative correction of the masks due to the higher computational demand of 3D segmentation. We run a grid search over the different options for deriving prompts, see Sec. 2.3, on separate validation data.
@@ 3 Experiments
We finetune different models based on SAM with ViT-b encoder, initialized with the weights from [33]: MedSAM, SimpleFT and MedicoSAM, see Sec. 2.2 for the respective training strategies. The models are trained on the publicly available subset [^3] The dataset contains a private test set of 920k images and 4 million masks. of SA-Med2D-20M [73], which contains 3.7 million images and 15.8 million masks. To compare the impact of $p_{box}$ and $n_{steps}$ (see Alg. 1), we train a model for each of the three configurations on 60% of the data (50% train, 10% val), using a stratified split over different modalities, setting $p_{mask}=0$, $e_{sem}=0$. We train another version of MedicoSAM on the full dataset (90% train, 10% val; same val split as before), with $p_{mask}=0.5$ and $e_{sem}=1$. The latter model is included to study the impact of masking and semantic decoder pretraining. The smaller subset for the three initial models is chosen to reduce training cost. The models are trained on 8 A100 GPUs with 80GB VRAM for 300,000 iterations with a batch size of 7 per GPU, corresponding to 21 epochs for models trained with 60% data, and 11 epochs for the model trained on the entire data. We use the AdamW optimizer [47] with an initial learning rate of 1e51e-5 and a scheduler that reduces the learning rate by a factor of 0.9 after each epoch. We evaluate these four models, the original SAM and SAM2 models, and published derived models for interactive segmentation (2D and 3D), semantic segmentation (2D and 3D), and integration with user-friendly tools.
Figure 3: a) Overall results for interactive 2D segmentation. We report the Dice coefficient for simulated interactive segmentation. Each bar corresponds to the result of a correction iteration, starting either from a point (green) or a box (yellow) prompt. The result after correction is highlighted in dark green / dark purple. We compare 10 different models. Models trained by us are marked with a * and the model trained on the entire dataset is marked in bold font. The same model notation is used in all figures. b) Interactive segmentation results for 16 individual datasets. We report the absolute difference of the Dice coefficient compared to the original SAM and report only the results for the initial and final segmentation.
@@ 3.1 Interactive Segmentation
We evaluate ten different models for interactive 2D segmentation. We use 16 datasets that are not part of our training dataset to evaluate generalization capabilities. These datasets represent a variety of medical segmentation tasks from CT [57, 3, 58], dermoscopy [14], endoscopy [64, 53], MRI [51, 16], ophthalmology [35], ultrasound [38, 48, 32], and X-Ray [29, 40, 7, 78]. Here, 3D dataset are split into separate images.
@@
The results averaged over all datasets are shown in Fig. 3 a). We report the results for initial prompt-based segmentation and correction iterations, see also Sec. 2.5, using a point or box as initial prompt. Fig. 3 b) shows the results for individual datasets, where we only report the results for the initial and final segmentation, corresponding to the last correction iteration. We report the difference in the Dice coefficient compared to SAM. MedicoSAM is the only model that clearly improves upon SAM in all settings. The other finetuned models either only improve for the segmentation with a single prompt (MedSAM*, SimpleFT*, MedSAM2) or lead to an overall worse segmentation. Models trained with $n_{steps}=0$ yield worse results for an increasing number of prompts. SAM2 performs slightly worse compared to (original) SAM.
Figure 4: Results for interactive 3D segmentation for 6 different datasets. We report the difference in Dice score compared to SAM for four other models. Segmentations are derived from a single point (green) or box (yellow) prompt placed in the central slice for each object in the respective dataset. We use the implementation of [1] for methods using SAM, determining the best method for prompt propagation on a separate validation set, see also Sec. 2.3. SAM2 supports 3D segmentation by default.
A qualitative comparison of results and image embeddings for different models is shown in Fig. 5. We also study the statistical reliability based on five random seeds on a single dataset in Fig. 6a, where we observe that the standard deviation of results is low compared to differences in the model performance. We perform an additional study to further understand the effect of $n_{steps}$ and $p_{mask}$ in Fig. 7.
@@
We evaluate interactive 3D segmentation with a single point or box derived from segmentation annotations for 6 different external datasets from MRI [56, 51], CT [23, 65], and ultrasound [32, 36]. We compare six different models. For the models based on SAM, we use the method described in Sec. 2.3 and we find the best setting with a grid search on a separate validation set. SAM2 supports interactive 3D segmentation as is (by interpreting the 3D data as a video), and we do not perform a grid search to optimize parameters for inference (same for MedSAM2). The overall results are shown in Fig. 1 c) (except for SimpleFT) with results for individual datasets in Fig. 4. Only MedicoSAM improves consistently. The fact that SAM2 and MedSAM2 are used as is, without optimizing parameters in a grid search, may disfavor them.
Figure 5: a) Qualitative results for interactive 2D segmentation. We compare interactive segmentation based on a single point or single box prompt (cyan) with SAM, MedSAM and MedicoSAM for nine different datasets. For each image, we show prompts with a large improvement of MedicoSAM over SAM and the corresponding MedSAM result. b) Outputs of the image encoder from the three different models on different datasets, additionally Abdominal CT [62] and Brain MRI [59] visualized by their three main PCA components. MedSAM and MedicoSAM seem to learn a more discriminative representation with clearer distinction of background.
Figure 6: a) Statistical analysis of interactive segmentation on the ABUS data for breast tumor segmentation in ultrasound images [2] with three models. We run each interactive segmentation experiment five times with different random seeds and report standard deviations as error bars. The deviations are small compared to differences in model performance. b) Statistical analysis of semantic segmentation (same data as a)). We run the training for each of the four models five times with different random seeds. The deviations are larger than performance differences, the model with pretrained decoder shows a lower deviation.) The methods are colored in decreasing shades of blue, darker indicates better results.
Figure 7: a) Ablation study for the influence of $n_{steps}$ in Alg. 1. We run interactive segmentation with 32 correction iterations, two with $n_{steps}=0$ (MedSAM, SimpleFT), two with $n_{steps}=8$. The former show a degrading performance with further prompts, the latter show increasing performance, but plateau after 6-8 iterations. b) Ablation study for the influence of $p_{mask}$, comparing models trained with $p_{mask}$ = 1, 0, 0.5 (SAM, MedicoSAM*, MedicoSAM*) and interactive segmentation without and with use of the previous mask prediction as prompt for the next iteration. SAM and MedicoSAM* perform slightly better in the setting corresponding to their training, MedicoSAM* is robust.
Figure 8: Semantic segmentation results for 2D and 3D segmentation. The names in italic font indicate 2D data and in bold font 3D data. We compare different pretrained models, original SAM, MedSAM, SimpleFT and MedicoSAM in three versions: trained on the reduced training set, trained on the entire training set (bold font) and trained on the entire training set and using decoder initialization (bold font, subscript ”Dec”). The three best SAM-derived methods are colored in decreasing shades of blue, darker indicates better results, gray otherwise. Results for reference methods are indicated by horizontal lines. The dotted horizontal line indicates a case where only a subset of classes could be segmented with BiomedParse.
@@ 3.2 Semantic Segmentation
We evaluate different pretrained SAM models for semantic segmentation in 2D and 3D, using the implementations described in Sec. 2.4. For 2D segmentation we use 6 external datasets from dermoscopy [10], mammography [40], narrow band imaging [64], optical coherence tomography [74], panoramic radiographs [63], and X-Ray angiography [7]. For 3D segmentation we use 6 external datasets from CT [65], MRI [51, 54, 56], and ultrasound [32, 36]. We use separate splits for training and evaluation. The results are shown in Fig. 8. We also report the results for nnU-Net [28] and Swin UNETR [19] trained on the same splits, using the default nnU-Net v2 setup and the MONAI [6] implementation, respectively. We also report the results for BiomedParse [81], a foundation model for text-based segmentation trained on a large biomedical imaging dataset. We apply it to the 2D datasets for which we could find a modality prompt matching the training data of BiomedParse.
Here, we see an advantage in using specifically pretrained backbones (MedSAM, SimpleFT, MedicoSAM) compared to the initial SAM model. We also see an advantage in using the pretrained segmentation decoder. These findings are consistent across 2D and 3D datasets. In 2D, SAM-derived models are competitive or slightly better than nnU-Net and better than Swin UNETR. In 3D, nnU-Net is the best method, in particular due to the poor performance of SAM-based models on two of the datasets. Note that nnU-Net and Swin UNETR are trained from scratch, whereas SAM-derived models are initialized with a pretrained encoder and in one case also a pretrained decoder. BiomedParse is applied to without further training and where applicable performs at-par with the best MedicoSAM model.
@@ 3.3 Tool Integration
An important practical aspect in improving SAM as a foundation model is the integration with tools for data annotation. Hence, models should not introduce changes to the architecture that make it incompatible with the SAM library and tools using it. We check this for three models, MedSAM, SAM-Med2D and MedicoSAM, and four different tools: two napari plugins [15, 1] and two 3D Slicer extensions [46, 75]. Tab. 1 shows the compatibility. We find that SAM-Med2D does not work in any of the tools because it uses adapters and changes the image input size. MedSAM and MedicoSAM work in all of the tools with at most small code changes. We qualitatively compare the models for data annotation with these tools in Fig. 9.
@@
Figure 9: Using SAM, MedSAM and MedicoSAM in four different tools for interactive segmentation. The top row shows the reference image and the object that is annotated. The four rows below show interactive segmentation results with the different tools. Data on the left hand side shows a section from a abdominal MRI scan [31], the right hand side shows a section from a cervical CT scan [37].
4 Conclusion and Discussion
We have comprehensively studied how to improve SAM [33] as a foundation model for medical images, by evaluating the impact of different finetuning objectives on interactive and semantic segmentation. We found that interactive segmentation improved clearly, compared to original SAM and SAM2 [61], but critically dependent on the choice of objective. For semantic segmentation, we found that domain specific pretraining also provides a benefit. However, the segmentation quality is only modestly better than nnU-Net [28] in 2D and worse in 3D; despite nnU-Net not being pretrained. We also argued that models based on SAM should adhere to the original architecture to enable integration with user-friendly tools. This is especially important due to their improvement in interactive segmentation, which relies on such tools for practical value. We have published our best model, MedicoSAM, and believe that it will be of great practical value for data annotation. Furthermore, we believe that our findings will also prove valuable to adapt recent and future foundation models, e.g. SAM2, to medical imaging.
Our main goal was the comparison of different objectives for finetuning SAM on large medical data. We found that it is crucial to use an objective that trains segmentation with box and mask prompts for multiple training iterations, i.e. $0<p_{box}<1$ and $n_{steps}>0$ (Alg. 1). Otherwise we observed a catastrophic forgetting-like effect, where the model did not yield accurate results for interactive segmentation with a point prompt or with multiple prompts, see results for MedSAM and FT-SAM in Fig. 3. This observation is especially important since MedSAM, the most cited SAM version for medical images, performs worse than the original SAM in many settings due to its choice of objective, which has been independently reproduced, e.g. by MedSAMix [72]. While not explicitly studying other values than $p_{box}=\{0.5,1\}$, we assume that values smaller than 0.5 would favor the performance with point prompts and vice versa, but would likely not yield to catastrophic forgetting. A value of 0 would likely result in a model that yields poor results in response to box prompts. We found that models trained with a value of $n_{steps}=8$ were robust to segmentation with a larger number of prompts, i.e. their performance did not decrease with increasing number of prompts as is the case for $n_{steps}=0$. However, their performance plateaus. See Fig. 7a. Training with a larger $n_{steps}$ value could potentially delay the onset of this plateau and increase its height, i.e. lead to better segmentation with more prompts. Finally, we also studied the effect of $p_{mask}$, where we found that models trained with $p_{mask}=0$ perform better without a mask prompt in interactive segmentation, models trained with $p_{mask}=1$ perform better with a mask prompt, and models trained with $p_{mask}=0.5$ are robust. See Fig. 7b. While this trend is expected, interestingly, the difference in performance is not very pronounced, i.e. a model trained with $p_{mask}=1$ does not learn to rely on the presence of a mask prompt and vice versa.
We also studied the effect on semantic segmentation by finetuning adapted architectures for specific segmentation tasks. Here, we found that SAM-derived models performed modestly better than SAM, but that the exact objective used for pretraining them on the large medical dataset did not have a big impact (similar results for MedSAM, MedicoSAM and SimpleFT). A further modest improvement could be gained by using a pretrained segmentation decoder. See results in Fig. 1 d,e and Fig. 4. Note that the best SAM-derived model performs only modestly better than nnU-Net [28] in 2D segmentation and worse than in 3D, despite the fact that nnU-Net was not pretrained. This highlights that the advantages of SAM-derived models for semantic segmentation in medical images are not yet clear, at least with our implementation. In this context, the finding of [27] that transformer-based segmentation architectures generally do not perform better than nnU-Net is also relevant. Further improvements would likely require updates of the adaptation strategy, especially an extension to 3D pretraining for volumetric segmentation.
Another important emergent trend are foundation models that can segment images based on text prompts. They are especially promising for semantic segmentation without the need for further training. We have included a comparison to one such model, BiomedParse [81] in selected experiments. There are further models, in many cases based on the CLIP [60] architecture, for medical imaging [82, 42, 45]. MedCLIP-SAM [34] combines CLIP and SAM for text based segmentation. Hence, a promising future avenue of research is the combination of such a model with our findings, to obtain a foundation model that can address both SAM-style interactive and text-based segmentation.
Other related research, e.g SAM-REF [76], has investigated the effect of late and early fusion of image features and prompts. SAM uses a late fusion approach. Early fusion can lead to better segmentation of fine-grained structures, while being less efficient because it couples image and prompt processing. Hybrid approaches can provide a good trade-off. This work could profit from our findings on finetuning objectives, as these are orthogonal to the feature alignment mechanism. Furthermore, recent work on parameter efficient finetuning of SAM for biomedical images has found that strategies like LoRA do not provide better segmentation quality, but can lead to more efficiency [66]. These findings likely directly apply to semantic segmentation finetuning in our case.